OpenStudy (anonymous):

3M+N1 M-6N=13 WHATS THE SOLUTION ?

OpenStudy (saifoo.khan):

Hi Korie.

OpenStudy (anonymous):

HI HOW R YOU ?

OpenStudy (saifoo.khan):

i cant get the ques.

OpenStudy (saifoo.khan):

A lil tired but fine! :D

OpenStudy (anonymous):

WHAT ONES ARE YOU GOOD AT IM STUCK ON ALOT OF THEM O YA AND THIS IS JUST A CHECK POINT AND NOT A TEST I JUST NEED A LITTLE HELP

OpenStudy (anonymous):

Y U SO TIRED ?

OpenStudy (saifoo.khan):

are these simultanious equations? coz czns came along @ my place for dinner and all.. & i'm awake from 8am..

OpenStudy (anonymous):

i am really tired and overwhelmed with having a bby and going to work man wew but I am srtong I can do it ...:)

OpenStudy (saifoo.khan):

Oh. Okay!

OpenStudy (saifoo.khan):

so are these simultaneous equations?

OpenStudy (anonymous):

what about (4-r-8)+4>5(r+3)-22

OpenStudy (llort):

m=6n+13 3(6n+13)+n=0 19n+39=0 n=-39/19

OpenStudy (saifoo.khan):

For the simultaniuos eqs, there solutions are.. x = 13/19 y = -39/19

OpenStudy (llort):

3m-39/19=0

OpenStudy (saifoo.khan):

(4-r-8)+4>5(r+3)-22 -r -4 + 4 > 5r + 15 -22 -r -5r > -7 -6r > -7 6r < 7 therefore: r < 7/6

OpenStudy (anonymous):

so llort is the answer 3 m -39 over 19 = 0 ?

OpenStudy (anonymous):

wow u r so smart even when your tired

OpenStudy (saifoo.khan):

this is the ANSWER: M = 13/19 N = -39/19

OpenStudy (saifoo.khan):

Lol, thnks for that.

OpenStudy (llort):

I am assuming, if I do not see an equal sign, I assume it is either equal to x or 0 depending on if there is variables

OpenStudy (llort):

I was letting you finish... I don't like doing everything for people.

OpenStudy (anonymous):

ok so saifoo is that r <7/6 is that a fraction 7/6

OpenStudy (saifoo.khan):

Yup, it is a Fraction.

OpenStudy (saifoo.khan):

i used my calc, so cant be wrong! :)

OpenStudy (anonymous):

ok ty

OpenStudy (saifoo.khan):

uw.

OpenStudy (anonymous):

a saifoo how would i WRITE THAT

OpenStudy (anonymous):

M = 13/19 N = -39/19

OpenStudy (saifoo.khan):

Oh Yeas!

OpenStudy (anonymous):

SO IT IS JUST LIKE THAT

OpenStudy (saifoo.khan):

Yo!

OpenStudy (anonymous):

WASUP

OpenStudy (saifoo.khan):

Chatting. Tired, Helping ppl here! :)

OpenStudy (anonymous):

THATS COOL

OpenStudy (anonymous):

SO DO I WRITE IT JUST LIKE YOU DID

OpenStudy (saifoo.khan):

Yeah!

OpenStudy (anonymous):

A can u meet me on twiddla i need help with graphing

OpenStudy (saifoo.khan):

Sure! :)

OpenStudy (saifoo.khan):

Link me!

OpenStudy (anonymous):

ok how do i do that

OpenStudy (saifoo.khan):

http://www.twiddla.com/573008

OpenStudy (saifoo.khan):

here u go!

OpenStudy (anonymous):

http://www.twiddla.com/569181